If you shoot a bullet with a velocity of 300 m/s out of a barrel that is 2 m off of the ground, how far will the bullet go before it hits the ground?
1
Expert's answer
2020-11-12T09:16:47-0500
"x=vt=v\sqrt{\frac{2h}{g}}\\x=300\sqrt{\frac{2(2)}{9.8}}=192\ m"